OpenStudy (anonymous):

Write a recursive formula for the sequence 8, 10, 12, 14, 16, ... . Then Determine the next term in the sequence

OpenStudy (anonymous):

answer choices :

OpenStudy (anonymous):

a_1 = 8 ; a_n = a_(n-1) + 2

OpenStudy (anonymous):

O.0 ...

OpenStudy (anonymous):

so choice 1 :D

OpenStudy (anonymous):

how did you know that a_1 was 8 ?

OpenStudy (anonymous):

because you start with 8

OpenStudy (anonymous):

a1 = the first number

OpenStudy (anonymous):

a1=8 and a2=8+2 so an=8+2(n-1)

OpenStudy (anonymous):

does that kinda make sense? :)

OpenStudy (anonymous):

ohh, yea that makes sense! thanks :)

OpenStudy (anonymous):

yw :)